OpenStudy (anonymous):

An offshore oil well is leaking oil and creating a circular oil slick. If the radius of the slick is growing at a rate of 6 miles/hour, find the rate at which the area is increasing when the radius is 7 miles. (The area of a disc of radius r is A = πr2.)

OpenStudy (anonymous):

Need to find dA/dt when dr/dt = 3.5 and r = 0.5. A(r) = π r^2 Differentiate d/dt [A(r)] = d/dt [π r^2] dA/dr = π 2r (dr/dt) dA/dr = π 2(0.5) (3.5) dA/dr = 3.5 π

OpenStudy (anonymous):

Refer to the Mathematica calculation attached.